OpenStudy (anonymous):

evaluate integral: ln(x + sqrt(x^2 -1)) dx

OpenStudy (anonymous):

\[\int\limits \ln (x+\sqrt{x^{2}-1}) dx\]

ganeshie8 (ganeshie8):

try by parts

OpenStudy (anonymous):

I know you let \[x = \sec^2 \theta\] \[dx = 2\sec^2 \theta \tan \theta d \theta\] so \[2 \int\limits \ln(\sec^2 \theta +\tan \theta ) (\sec^2 \theta \tan \theta) d \theta\]

ganeshie8 (ganeshie8):

\[\int \ln \spadesuit~dx = x\ln \spadesuit - \int x(\ln \spadesuit )' ~dx \]

ganeshie8 (ganeshie8):

\[\int \ln (x+\sqrt{x^{2}-1}) dx =x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\ln (x+\sqrt{x^{2}-1})\right)' dx \]

ganeshie8 (ganeshie8):

oh did they cover integration by parts yet ?

OpenStudy (anonymous):

we went over integration by parts already, we're on the part where they combined all integration strategies, and we're trying to familiarize which one we use when we see a problem.

ganeshie8 (ganeshie8):

good, integration by parts is best for this problem give it a try

ganeshie8 (ganeshie8):

you should think of by parts if you see log term in integrand

OpenStudy (anonymous):

did you already start integration by parts for the equation you gave me? uv - ingetral vdu?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

this one is a hard one. did you let u = ln dx and dv = (x + sqrt(x^2-1) ?

ganeshie8 (ganeshie8):

\( u = \ln (x+\sqrt{x^{2}-1}) \) \(\large v = x\)

ganeshie8 (ganeshie8):

\[\large (uv)' = uv' + u'v \implies uv = \int uv' + u'v\]

OpenStudy (anonymous):

i was taught to let something be u and other be dv. you derive u to get du, and integral dv to get v.

ganeshie8 (ganeshie8):

yes, dv = 1 you get v = x

OpenStudy (anonymous):

ah ha....

OpenStudy (anonymous):

is there a reason why you let the whole thing equal u. is there something that pops out that made you do that?

OpenStudy (anonymous):

or is it trial and error and as you do more practice problem you know what to do.

ganeshie8 (ganeshie8):

letting that whole thing equal to u comes naturally once you know how by parts works

ganeshie8 (ganeshie8):

\[\large \int \ln (x+\sqrt{x^{2}-1}) dx = \int \ln (x+\sqrt{x^{2}-1}) \color{red}{\cdot 1}~ dx\]

ganeshie8 (ganeshie8):

the antiderivative of 1 is x so you want to integrate 1, doing anything else is ridiculolus here

OpenStudy (anonymous):

okay, i got the equation that you posted earlier.

OpenStudy (anonymous):

i was just trying to understand why you decided to let u be = to that and dv = to 1

ganeshie8 (ganeshie8):

thats a good question actually :) heard of ILATE mnemonic for by parts ?

OpenStudy (anonymous):

no, i have not heard of that.

ganeshie8 (ganeshie8):

that helps u in deciding what to pick for dv fast

ganeshie8 (ganeshie8):

nope. it helps u in deciding what to pick for `u` fast

OpenStudy (anonymous):

so i is first, L is second, A is third etc etc?

OpenStudy (anonymous):

oh you let exponential go first, then T, then A?

ganeshie8 (ganeshie8):

yes if u see a `log` term and an `exponent` in your integrand, you pick `log` as `u`

ganeshie8 (ganeshie8):

exponential is always the last preference for `u` because it wont reduce no matter how many times you differentiate

OpenStudy (anonymous):

ahhh, okay. let me see what i can do with the resource you gave me.

OpenStudy (anonymous):

algebra are x^2 x^3, and exponentials are e^x right?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

For your present problem : \[\begin{align}&\int \ln (x+\sqrt{x^{2}-1}) dx =x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\ln (x+\sqrt{x^{2}-1})\right)' dx\\~\\&= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(x+\sqrt{x^{2}-1}\right)' dx\end{align}\]

OpenStudy (anonymous):

is the answer \[= xln(x+\sqrt{x^2-1}) - arcsec(x) + C\]

OpenStudy (anonymous):

i don't think i derived u correctly.

ganeshie8 (ganeshie8):

\[\begin{align}&\int \ln (x+\sqrt{x^{2}-1}) dx =x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\ln (x+\sqrt{x^{2}-1})\right)' dx\\~\\&= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(x+\sqrt{x^{2}-1}\right)' dx \\~\\ &= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(1+\dfrac{2x}{2\sqrt{x^{2}-1}}\right) dx \end{align}\]

OpenStudy (anonymous):

okay i got that.

ganeshie8 (ganeshie8):

\[\begin{align}&\int \ln (x+\sqrt{x^{2}-1}) dx =x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\ln (x+\sqrt{x^{2}-1})\right)' dx\\~\\&= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(x+\sqrt{x^{2}-1}\right)' dx \\~\\ &= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(1+\dfrac{2x}{2\sqrt{x^{2}-1}}\right) dx\\~\\ &= x\ln (x+\sqrt{x^{2}-1}) - \int x \left(\dfrac{1}{x+\sqrt{x^{2}-1}} \right)\left(\dfrac{\sqrt{x^{2}-1}+x}{\sqrt{x^{2}-1}}\right) dx\\~\\ &= x\ln (x+\sqrt{x^{2}-1}) - \int\left( \dfrac{x}{\sqrt{x^{2}-1}}\right) dx\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

substitute `u^2 = x^2-1` to evaluate the integral

OpenStudy (anonymous):

is there a reason why you did u^2 = x^2 -1 instead of u = x^2 -1

OpenStudy (anonymous):

is it because its under a sq rt?

zepdrix (zepdrix):

Yooooon! Your original approach was just fine. You were just off by a little bit with your substitution. You wanted: \(\Large\rm x=\sec\theta\)\[\Large\rm dx=\sec\theta\tan\theta~d\theta\] Then,\[\large\rm \int\limits \ln\left(x+\sqrt{x^2-1}\right)dx=\int\limits \ln\left(\sec \theta+\tan \theta\right)\sec \theta \tan \theta~d \theta\] And you can do parts from here :) \[\Large\rm u=\ln(\sec \theta+\tan \theta), \qquad \qquad dv=\sec \theta \tan \theta~d \theta\] The other approach seems good also though. ^^ This way is maybe a bit easier to mess up.

OpenStudy (anonymous):

for \[\int\limits \frac{ x }{ \sqrt{x^{2}-1} } dx\] i got \[\frac{ du }{ 2\sqrt{u} }\] is that correct?

zepdrix (zepdrix):

You should get that BEFORE integrating. Is that what you meant?

zepdrix (zepdrix):

Oh you have a du in there, so yah that would make sense :) lol Integrate! >.<

OpenStudy (anonymous):

is the integration 4u^(1/2)?

OpenStudy (anonymous):

ahhh, i made a mistake its u^(1/2) okay i got the answer, thank you all for helping out!!!